Ashley is 10 years older than her sister. In six years she will be twice her sister's age. How old is Ashley?
Answer by TimothyLamb(4379)

a = s + 10
a + 6 = 2(s + 6)
---
a + 6 = 2s + 12
(s + 10) + 6 = 2s + 12
s + 10 + 6 = 2s + 12
---
Answer:
s = 4
a = 14
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php